A wire 1000 ft long has a resistance of 0.5 kΩ and an area of 94 CM. Of what material is the wire made (T= 20°C)?
1
Expert's answer
2020-03-05T07:58:35-0500

Resistivity of various materials:

"\rho=R\frac{A}{l}"

A - area in circular mils (CM)

l - length (feet)

"R=0.5k\Omega=500\Omega - resistance"

"\rho=500\frac{94CM}{1000ft}=47 ~CM\cdot \Omega/ft ~at~ 20°C"

the resulting resistivity value matches Nickel

Answer: Nickel